OpenStudy (ashking1):

What is the simplified form of sqrt128n^17

OpenStudy (anonymous):

ur sqrt encompasses wat values? is it sqrt(128) n^17? or sqrt(128 n^17)

OpenStudy (ashking1):

\[\sqrt{128n ^{17}}\]

OpenStudy (anonymous):

\[\sqrt{128n ^{17}} = \sqrt{128} \times \sqrt{n^{17}}\] \[= \sqrt{2 \times 4 \times 4 \times 4} \times n^{\frac{17}{2}}\] \[= 2 \times 2 \times 2 \times \sqrt{2} \times n^{\frac{17}{2}}\] \[= 8\sqrt{2}n^{\frac{17}{2}}\]